# AEGIS : Automated Co-Evolutionary Framework for Guarding Prompt Injections Schema

### v1

@@ Abstract
Prompt injection attacks pose a significant challenge to the safe deployment of Large Language Models (LLMs) in real-world applications. While prompt-based detection offers a lightweight and interpretable defense strategy, its effectiveness has been hindered by the need for manual prompt engineering. To address this issue, we propose AEGIS , an Automated co-Evolutionary framework for Guarding prompt Injections Schema. Both attack and defense prompts are iteratively optimized against each other using a gradient-like natural language prompt optimization technique. This framework enables both attackers and defenders to autonomously evolve via a Textual Gradient Optimization (TGO) module, leveraging feedback from an LLM-guided evaluation loop. We evaluate our system on a real-world assignment grading dataset of prompt injection attacks and demonstrate that our method consistently outperforms existing baselines, achieving superior robustness in both attack success and detection. Specifically, the attack success rate (ASR) reaches 1.0, representing an improvement of 0.26 over the baseline. For detection, the true positive rate (TPR) improves by 0.23 compared to the previous best work, reaching 0.84, and the true negative rate (TNR) remains comparable at 0.89. Ablation studies confirm the importance of co-evolution, gradient buffering, and multi-objective optimization. We also confirm that this framework is effective in different LLMs. Our results highlight the promise of adversarial training as a scalable and effective approach for guarding prompt injections.
Introduction
@@
Large Language Models (LLMs) have rapidly become integral components of modern AI systems, powering a wide range of downstream applications such as education, healthcare, and software engineering. However, the deployment of LLMs in real-world settings also exposes them to security risks, most notably prompt injection attacks, where maliciously crafted inputs manipulate the model into producing unintended or harmful outputs. Unlike traditional adversarial examples, prompt injections exploit the semantic and contextual flexibility of natural language, making them particularly challenging to detect and defend against.
@@
In this work, we introduce AEGIS , a novel adversarial co-evolution framework for automated discovery of robust defenses against prompt injection attacks. Our framework jointly evolves attack and defense prompts in an iterative, GAN-inspired process, where attackers continuously refine adversarial strategies and defenders adapt in response. At the core of our method is TGO+, an enhanced textual gradient optimization module that simulates gradient-like updates using natural language feedback. By leveraging multi-route optimization signals, a gradient buffer, and adversarial co-training, AEGIS autonomously explores the space of defensive strategies without requiring model fine-tuning or human-crafted rules.
We evaluate our framework in the context of automated assignment grading, a realistic scenario where malicious prompts can manipulate grading outcomes. Across multiple LLMs, AEGIS consistently improves both attack and defense prompts over successive iterations, achieving state-of-the-art robustness against real-world injection attempts while preserving utility on benign inputs. Furthermore, ablation studies confirm the critical role of co-evolution and gradient replay in sustaining long-term robustness.
Contributions
Our main contributions are as follows:
  • We propose a general co-evolutionary adversarial framework that systematically evolves both attackers and defenders, enabling adaptive robustness against prompt injection attacks.
  • We design TGO+, an enhanced prompt optimization method with multi-route textual gradients and gradient buffering, tailored for black-box LLMs.
  • We demonstrate the effectiveness of our framework in a specific real-world application, conducting comprehensive experiments on authentic datasets and multiple LLMs to showcase superior defense performance and strong cross-model generalizability.
@@
Related Works
Under the scenario of applying LLMs as black-box components in downstream applications, our work focuses on developing an automatic framework for discovering reliable defenses against prompt injection attacks—without relying on any additional training resources or model fine-tuning.
@@
A key observation we make is that most existing training-free defenses against prompt injection still depend heavily on human-crafted design or heuristic insights. In this section, we highlight three recent yet fundamentally different approaches, each representative of a popular defense strategy, and show how they align with our observation. This motivates the need for a more general and automated framework such as ours.
Training-Free Defenses against Prompt Injections
Several recent training-free defenses rely on manually defined prompt formulations or input transformation schemes, which limits their robustness and adaptability. We briefly introduce three representative methods:
@@
  • LLM-based Detection Prompts. PromptArmor (Shi et al. 2025) turns an off-the-shelf LLM into a detector by prompting it to identify and remove injected content. However, it depends on a manually crafted detection prompt, which is fixed and not optimized for different threat settings.
@@
  • Input-level Structure Encoding. Spotlighting (Hines et al. 2024) inserts manually chosen provenance markers to distinguish user input from system instructions. The choice of marker syntax and format is hand-engineered and remains static across tasks.
@@
  • Behavioral Consistency Checking. MELON (Zhu et al. 2025) detects indirect prompt injections by re-executing the task with a fixed masking prompt. Although task-agnostic and effective, its masking strategy is manually specified and may not generalize across diverse attacker behaviors.
@@
All three methods share a common limitation: their reliance on fixed, manually designed prompts or transformation templates. In contrast, our framework automates the exploration of the prompt and transformation space, enabling adaptive and potentially more robust defenses against a broader range of injection attacks.
Prompt Optimization
In addition to demonstrating how our framework can enhance the search process for robust training-free defenses, we also position our work within the broader landscape of prompt optimization methodologies. Following the terminology used in Cui et al. (2025), our work introduces a novel adversarial prompt optimization framework that falls under the category of heuristic-based prompt search algorithms. Cui et al. (2025) categorize prompt optimization techniques by their target objectives, including task performance, generalizability, safety and ethical constraints, and multi-objective optimization.
To facilitate comparison, we further organize existing objectives into two broad categories, based on whether the optimization setting assumes a fixed or evolving task environment. Although both directions have important practical applications, the evolving task environment that requires interactive objective optimization setting is inherently more challenging due to its dynamic and adversarial nature. Moreover, this setting more accurately reflects the core difficulty of many prompt-based defense methods against malicious user behavior in real-world LLM applications.
Static Objective Prompt Optimization.
@@
This category includes works that optimize prompts for known and fixed objectives, where the task definition and evaluation criteria remain constant throughout. Static Objective Optimization is therefore suited for fixed task environments and does not account for evolving threats or task definitions.
@@
  • Task-Specific Optimization. These approaches aim to improve performance on specific downstream tasks under static conditions, optimizing metrics such as accuracy or BLEU without addressing adversarial dynamics. Pryzant et al. (2023) propose ProTeGi, a heuristic method that refines prompts using natural language feedback called textual gradients, beam search, and bandit-based selection. Chen et al. (2024) introduce PROMST, which integrates rule-based and learned heuristics to optimize instructions and demonstrations across multi-step tasks. Opsahl-Ong et al. (2024) present MIPRO, a black-box meta-optimization framework that jointly refines prompts in multi-stage LM programs using program-aware proposals and surrogate evaluation.
@@
  • Cross-Domain Generalization. Li et al. (2024) propose Concentrate Attention, which uses attention strength and stability in deeper transformer layers to guide prompt optimization. They introduce a concentration-based loss for soft prompts and a reinforcement learning strategy for hard prompts, achieving better out-of-domain performance without sacrificing in-domain accuracy.
@@
  • Static Multi-Objective Optimization. Sinha et al. (2024) propose Survival of the Safest (SoS), a multi-objective evolutionary framework that optimizes prompts for both task performance and safety. By interleaving semantic and feedback-based prompt mutations, SoS identifies candidate prompts that balance accuracy and robustness, under a fixed threat model.
Interactive Objective Prompt Optimization.
@@
This category includes works that adapt to evolving task demands or adversarial settings. It is suited for evolving task environments, where objectives or threats may evolve over time. Despite its practical relevance, this direction has received relatively limited attention, likely due to the complexity of modeling dynamic behaviors and adversaries. A key real-world motivation for this setting is developing prompt-based defenses against malicious user behavior in LLM applications, such as prompt injection or jailbreak attacks. These scenarios demand optimization frameworks capable of anticipating and resisting adaptive adversarial strategies.
Robust Prompt Optimization (RPO), proposed by (Zhou, Li, and Wang 2024), is closely related to our work and shares a similar optimization objective. However, beyond its white-box setting, RPO and our approach also diverge in more fundamental ways—particularly in the underlying threat models dictated by the targeted scenarios. Specifically, RPO is designed to address jail-breaking attacks, whereas our work focuses on prompt injection attacks. This distinction leads to important differences in the problem formulation, as outlined below.
  • Attacker’s perspective: Unlike in the jailbreaking scenario, prompt injection attacks typically occur in application settings where the attacker is an end-user and thus subject to strict query limits imposed by the application owner. These constraints make it impractical to use gradient-based optimization or other approaches that require a large number of model queries. As a result, attackers often rely on semantic search strategies, reusing or adapting known effective attack prompts rather than optimizing them through extensive trial and error.
  • Defender’s perspective: In contrast to jailbreak detection, where the model’s misbehavior is usually explicit and easier to flag, prompt injection defenses are more nuanced. Here, the defender often has more control over the prompt construction pipeline, including both the user input and the system instructions, before sending the full prompt to the LLM. However, despite the stronger control, detecting prompt injections can be far more ambiguous. For example, in a grading application where an LLM scores both a high-quality and a low-quality article with 10/10, an attack may only be present in the low-quality article if it includes hidden or indirect instructions influencing the grading.
These differences in attacker capabilities and defense challenges result in fundamentally different threat assumptions and motivate distinct architectural and algorithmic choices in the design of our framework compared to RPO.
Method
@@
We propose a general adversarial co-evolution framework that enables both attackers and defenders to evolve automatically through iterative optimization. While our method can be applied to various scenarios, we demonstrate its effectiveness in the context of automated assignment grading. In this setting, an attacker attempts to obtain high scores by injecting adversarial prompts, and the defender aims to prevent misgrading. By leveraging LLM-guided feedback and prompt refinement, our system continuously improves both attack and defense strategies without human intervention.
@@
The core training procedure illustrated in Figure 2. In this framework, the attacker and defender evolve in alternating turns. In each cycle, the attacker evolves for a fixed number of iterations based on the current best defense. Once the attacker is fixed, the defender then evolves in response to the best attack from the previous step. This process continues until both sides converge (i.e., no further improvements) or a predefined maximum number of GAN iterations is reached. We describe different modules in the following sections.

### v2

@@ Abstract
Prompt injection attacks pose a significant challenge to the safe deployment of Large Language Models (LLMs) in real-world applications. While prompt-based detection offers a lightweight and interpretable defense strategy, its effectiveness has been hindered by the need for manual prompt engineering. To address this issue, we propose AEGIS , an Automated co-Evolutionary framework for Guarding prompt Injections Schema. Our method employs a two-level optimization process: at the inner loop, we leverage existing prompt optimization frameworks to refine individual prompts, while at the outer loop, adversarial agents exchange feedback to co-evolve and improve beyond standalone optimization. We then evaluate our system on a real-world assignment grading dataset of prompt injection attacks and demonstrate that our method consistently outperforms existing baselines, achieving superior robustness in malicious prompt detection. In particular, our defense improves the true positive rate (TPR) by 0.20 compared to the previous state of the art, with only a slight decrease in the true negative rate (TNR) of 0.02. Ablation studies confirm the importance of co-evolution, gradient buffering, and multi-objective optimization. We also confirm that this framework is effective in various LLMs. Our results highlight the promise of adversarial training as a scalable and effective approach for guarding against prompt injection attacks.
AEGIS : Automated Co-Evolutionary Framework for Guarding Prompt Injection
Ting-Chun Liu   Ching-Yu Hsu   Kuan-Yi Lee   Chi-An Fu   Hung-yi Lee Electrical Engineering, National Taiwan University {b10901039,b10901036,b10901091,b11901174}@ntu.edu.twhungyilee@ntu.edu.tw
[…]
1 Introduction
@@
Large Language Models (LLMs) have rapidly become integral components of modern AI systems, powering a wide range of downstream applications such as education. However, the deployment of LLMs in real-world settings also exposes them to security risks, most notably prompt injection attacks, where maliciously crafted inputs manipulate the model into producing unintended or harmful outputs. Unlike traditional adversarial examples, prompt injections exploit the semantic and contextual flexibility of natural language, making them particularly challenging to detect and defend.
@@
In this work, we introduce AEGIS , a novel adversarial co-evolution framework for automated discovery of robust defenses against prompt injection attacks, as illustrated in Figure 1. Our framework jointly evolves attack and defense prompts in an iterative, GAN-inspired process, where attackers continuously refine adversarial strategies and defenders adapt in response. The core of AEGIS is TGO+, an enhanced textual gradient optimization module that simulates gradient-like updates using natural language feedback. By leveraging multi-route optimization signals, a gradient buffer, and adversarial co-training, AEGIS autonomously explores the space of defensive strategies without requiring model fine-tuning or human-crafted rules.
We evaluate our framework on automated assignment grading, a realistic scenario where malicious prompts can manipulate grading outcomes. Across multiple LLMs, AEGIS consistently improves both attack and defense prompts over successive iterations, achieving state-of-the-art robustness against real-world injection attempts while preserving utility on benign inputs. Furthermore, ablation studies confirm the critical role of co-evolution and gradient replay in sustaining long-term robustness.
Our main contributions are as follow
  • Proposing a general co-evolutionary adversarial framework that systematically evolves both attackers and defenders, enabling adaptive robustness against prompt injection attacks.
  • Designing TGO+, an enhanced prompt optimization method with multi-route textual gradients and gradient buffering, tailored for black-box LLMs.
  • Demonstrating the effectiveness of our framework in a specific real-world application, conducting comprehensive experiments on authentic datasets and multiple LLMs to showcase superior defense performance and strong cross-model generalizability.
@@
2 Related Works
@@
A key observation we make is that most existing training-free defenses against prompt injection still depend heavily on human-crafted design or heuristic insights. In this section, we highlight three recent yet fundamentally different approaches, each representative of a popular defense strategy, and show how they align with our observation.
2.1 Training-Free Defenses
@@
  • LLM-based Detection Prompts. PromptArmor Shi et al. (2025) prompts an LLM to identify and remove injected content, but the detection prompt is manually crafted and not optimized for varying threat settings.
@@
  • Input-level Structure Encoding. Spotlighting Hines et al. (2024) inserts provenance markers to separate user input from system instructions, though the marker format is hand-engineered and static across tasks.
@@
  • Behavioral Consistency Checking. MELON Zhu et al. (2025) re-executes tasks with a fixed masking prompt to detect indirect injections. While task-agnostic, its masking strategy is still manually specified and may not cover diverse attacker behaviors.
@@
All three methods rely on fixed, manually designed prompts or templates. Our framework instead automates prompt and transformation exploration, enabling adaptive, more robust defenses against diverse injection attacks.
2.2 Prompt Optimization
Our work connects to the broader landscape of prompt optimization. Following the terminology used in Cui et al. (2025), our prompt optimization framework falls under the category of heuristic-based prompt search algorithms. Cui et al. (2025) categorize prompt optimization techniques by their target objectives. To facilitate comparison, we further organize existing objectives into two broad categories, based on whether the optimization setting assumes a static or evolving task environment.
2.2.1 Under Static Environment
@@
This category includes works that optimize prompts for known and fixed objectives, where the task definition and evaluation criteria remain constant throughout.
@@
Task-Specific Optimization aim to improve performance on specific downstream tasks under static conditions, optimizing metrics such as accuracy or BLEU without addressing adversarial dynamics. Pryzant et al. (2023) propose ProTeGi, a heuristic method that refines prompts using natural language feedback called textual gradients, beam search, and bandit-based selection. Chen et al. (2024) introduce PROMST, which integrates rule-based and learned heuristics to optimize instructions and demonstrations across multi-step tasks. Opsahl-Ong et al. (2024) present MIPRO, a black-box meta-optimization framework that jointly refines prompts in multi-stage LM programs using program-aware proposals and surrogate evaluation.
@@
Cross-Domain Optimization, for example, Li et al. (2024) propose Concentrate Attention, which uses attention strength and stability in deeper transformer layers to guide prompt optimization. They introduce a concentration-based loss for soft prompts and a reinforcement learning strategy for hard prompts, achieving better out-of-domain performance without sacrificing in-domain accuracy.
@@
Multi-Objective Optimization, such as Sinha et al. (2024), propose Survival of the Safest (SoS), a multi-objective evolutionary framework that optimizes prompts for both task performance and safety. By interleaving semantic and feedback-based prompt mutations, SoS identifies candidate prompts that balance accuracy and robustness, under a fixed threat model.
2.2.2 Under Evolving Environment
@@
This category includes works that adapt to evolving task demands or adversarial settings. It is suited for evolving task environments, where objectives or threats may evolve over time. This direction has received relatively limited attention due to the complexity of modeling dynamic behaviors.
Robust Prompt Optimization (RPO), proposed by Zhou et al. (2024), shares a similar adversarial optimization framework with ours. However, it adopts a white-box setting, where the defender simulates the attacker’s optimization process using gradient-based methods (e.g., GCG) on open-source models. For closed-source models, the defensive suffixes are optimized on open-source surrogates and then directly transferred for evaluation without further adaptation on the black-box target. In contrast, our method assumes a black-box setting, where attacker and defender are optimized without access to model internals or gradients.
3 Method
@@
We propose a general adversarial co-evolution framework that enables both attackers and defenders to evolve automatically through iterative optimization. Although the framework can be applied to different security-sensitive tasks, we focus on automated assignment grading as a concrete scenario to demonstrate its effectiveness. In this setting, an attacker attempts to obtain high scores by injecting adversarial prompts, and the defender aims to prevent misgrading. By leveraging LLM-guided feedback and prompt refinement, our system continuously improves both attack and defense strategies without human intervention.
@@
The core training procedure is illustrated in Figure 2. In this framework, the attacker and defender evolve in alternating turns. In each cycle, the attacker evolves for a fixed number of iterations based on the current best defense. Once the attacker finishes its training and gets the current best attack, the defender will evolve in response to it. This process continues until both sides converge (i.e., no further improvements) or a predefined maximum number of GAN iterations is reached. The overall algorithm can be seen in Appendix A.1.
We describe different modules in the following sections.
3.1 Attacker and Defenser
3.1.1 Attacker
The attacker module aims to generate adversarial prompts that will be further injected to the original prompt and can mislead the system to give a higher score. Within the co-evolutionary framework, the attacker evolves in alternating turns against a fixed defender, simulating an arms race between offensive and defensive strategies.
During each attacker iteration, a training phase is initiated to explore new adversarial candidates. Specifically, new attack candidates ($ATK\_Cand_{j}^{i}$) are generated using the $TGO^{+}$ module, which synthesizes gradient-like signals derived from grading feedback and guides the editing of existing prompts to enhance their adversarial strength. The attacker will then maintain a top-$k$ pool of the strongest attack prompts from previous cycles, denoted as $ATK^{i}_{j}$, by evaluating their effectiveness against the current best defense using $Eval()$ in the grading system.
To evaluate the effectiveness of each new attack candidate, we ask the LLM in the grading system to output three important values based on the original input and the attack candidate:
  • $S_{\text{benign}}$: The score assigned to the original (benign) input without any adversarial prompts.
  • $S_{\text{attacked}}$: The score assigned after injecting adversarial prompts into the original input.
  • ASR: The attack success rate, which quantifies the probability that an attack bypasses detection by the defense.
After generating these values, we first calculate the relative score change $\Delta S_{\text{rel}}$ in Equation (1):
where $S_{\text{benign}}$ means the maximum possible score defined by the grading system.
If $S_{\text{benign}}<S_{\text{max}}$, the relative score change is computed as the ratio between the observed improvement due to the attack and the maximum improvement that could possibly be achieved. Conversely, if $S_{\text{benign}}\geq S_{\text{max}}$, no further improvement is feasible, and $\Delta S_{\text{rel}}$ is set to zero. This normalization is critical because it accounts for the diminishing significance of score increments near the upper end of the grading scale. For instance, an increase from 8 to 9 carries greater weight than an increase from 1 to 2, as improvements become progressively harder to obtain as scores approach the maximum.
After computing the relative score change $\Delta S_{\text{rel}}$, we rank and maintain the top-$k$ adversarial attacks using the attack score defined in Equation (2).
$S_{\text{attack}}=w_{\text{asr}}\cdot(\text{ASR})^{p_{\text{asr}}}+w_{\text{sc}}\cdot(\Delta S_{\text{rel}})^{p_{\text{sc}}}$ (2)
where $w_{asr}$ and $w_{sc}$ are weights for the ASR and $\Delta S_{rel}$, respectively, and $p_{asr}$ and $p_{sc}$ are power parameters to control the sensitivity of each term.
The weighting coefficients ($w$) determine the relative importance of the two metrics in the overall score—for example, assigning a larger $w_{\text{asr}}$ prioritizes attacks that are more difficult for the defense to detect. The power parameters ($p$) modulate how strongly changes at different regions of the metric’s range influence the score—for instance, a larger $p_{\text{asr}}$ amplifies the effect of improvements in high-success regions (e.g., from 0.8 to 0.9) relative to low-success regions (e.g., from 0.1 to 0.2).
After the training procedure, the best-performing attack ($ATK^{i}_{best}$) from the top-$k$ pool is chosen with the highest attack score on the validation set using $Val()$ in the grading system. Then, the selected adversarial prompts will be fixed for the next defender evolution cycle, ensuring that the defender is trained against the most challenging known threat at the time.
3.1.2 Defender
Following the co-evolutionary setup, the defender evolves in response to a fixed attacker. After each attacker cycle, the best-performing attack prompt is used as the evaluation context against which the defender is trained. The evolution process mirrors that of the attacker: new defense candidates ($DEF\_Cand^{i}_{j}$) are generated via the $TGO^{+}$ module, and a top-$k$ pool of defense prompts denoted as $DEF^{i}_{j}$ is maintained via $Eval()$.
To evaluate the effectiveness of the defense prompt, we ask the LLM in the grading system to output two important values: True Positive Rate (TPR), which denotes the probability that the defense correctly detects an attack, and True Negative Rate (TNR), which denotes the probability that the defense correctly identifies a benign input as non-attacked.
After generating these two values, we will calculate the defense score in Equation (3):
$S_{defense}=w_{tp}\cdot(TPR)^{p_{tp}}+w_{tn}\cdot(TNR)^{p_{tn}}$ (3)
where $w_{tp}$ and $w_{tn}$ are weights for the True Positive Rate (TPR) and True Negative Rate (TNR), respectively, and $p_{tp}$ and $p_{tn}$ are their corresponding power parameters. Similar to the attack score, the coefficients and power parameters in the defense score allow for a nuanced and flexible evaluation of the effectiveness of the defense prompt.
Similarily, the best defense ($DEF^{i}_{best}$) is chosen from the top-$k$ pool with the highest defense score on the validation set using $Val()$ in the grading system after the training process.
3.2 TGO+ for Prompt Optimization
Inspired by the TGO framework proposed in  Pryzant et al. (2023), we adopt a modular design to optimize prompts through gradient-like updates in natural language space. Each TGO+ module is dedicated to a specific optimization goal (e.g., attack or defense) and operates in two stages: gradient acquisition and gradient application, as illustrated in Figure 3. The overall algorithm for TGO+ can be seen in Algorithm 2.
3.2.1 Gradient Acquisition
For each prompt and its gradient results (ASR, $\Delta S_{\text{rel}}$ for attack prompt; TPR, TNR for defense prompt), the module first collects the errors on these gradient results. These errors are then combined with a task-specific instruction and prompted to LLM, and LLM returns several feedback messages serving as the textual gradients—i.e., suggestions indicating how the prompt could be improved.
To ensure diverse learning, recent feedback messages are stored in a gradient buffer. This buffer encourages diversity in the optimization trajectory by prompting the LLM to generate alternative gradients even when the same input prompt is provided.
3.2.2 Gradient Application
In the second stage, the feedback is synthesized into a set of guidance messages based on the optimization type (e.g., ASR optimization for attack / TPR optimization for defense). These are used to construct an edit prompt that instructs the LLM to revise the original candidate prompt accordingly. After all these prompt candidates are generated, they are sent to $Eval()$ in the grading system to evaluate their effectiveness.
TGO+ enables gradient-like prompt updates without requiring access to model internals, making it compatible with black-box LLMs such as GPT-4o and Gemini-2.5-flash.
3.2.3 Key Innovations
  • Multi-Route Gradient Optimization: To enhance the optimization process, we employ a multi-route gradient strategy. This means that for each prompt, we generate textual gradients based on multiple optimization type. For instance, the Attackerś prompts are optimized based on either ASR or the relative score change. Similarly, the Defenderś prompts are optimized based on TPR or TNR. This allows for a more holistic and effective optimization process.
4 Experimental Setup
4.1 Dataset
4.2 Experimental Procedure
4.3 Hyperparameters
The default hyperparameters used in our experiments are summarized in Appendix A.3. These settings were used for the baseline experiment, and variations are explored in the ablation studies.
5 Results
5.1 Overall Evaluation
We evaluate the defense effectiveness of our framework against three baseline mechanisms, including Perplexity-based Detection Alon and Kamfonas (2023), LLaMA 3.1 Guard Inan et al. (2023), and the "Human-Crafted Prompt" defense, which is proposed in Chiang et al. (2024) to defend against real-world attacks (See Appendix A.6). For all these methods, we evalaute the defense against the real-world articles (147 malicious articles + 100 benign articles). The results, summarized in Table 1, show that our method achieves state-of-the-art defense performance. We present our results at both an early stage (Iteration 4) and the final stage (Iteration 8) of the GAN training.
Defense Method TPR TNR Human-Crafted Prompt Chiang et al. (2024) 0.64 0.91 Perplexity-based Detection Alon and Kamfonas (2023) 0.54 0.73 LLaMA 3.1 Guard Inan et al. (2023) 0.61 0.81 Our Method (AEGIS @ Iteration 4) 0.76 0.88 Our Method (AEGIS @ Iteration 8) 0.84 0.89
Table 1: Defense Method Comparison. The result shows that our method already achieves the best TPR at iteration 4, despite a slight decrease in TNR compared with Human-Crafted Prompt. The model even performs better at iteration 8, with better TPR and TNR compared with our method at iteration 4.
5.2 Iterative Performance
To illustrate the co-evolution of the attacker and defender, Table 2 shows the real-world evaluation for both agents at different iteration of the GAN training process. To be more specific, we evaluate the generated attacks against the defense in Human-Crafted Prompt from Chiang et al. (2024), and evaluate the generated defenses against the real-world articles (143 malicious articles + 100 benign articles) to calculate the TPR and TNR. The results demonstrate a clear trend of mutual improvement, where each agent becomes progressively stronger by adapting to the other.
The trend shows the attacker’s metrics (ASR, Relative Score Change) steadily increasing as it learns to bypass the improving defender. Simultaneously, the defender’s TPR improves as it learns to find the defense prompt to detect the adversarial attacks. This dynamic demonstrates a successful adversarial training loop, leading to highly robust agents.
Appendix A.4 provides examples of real prompt refinements, illustrating that the defense yields meaningful qualitative improvements.
5.3 Cross-Model Generalizability
To evaluate the robustness and generalizability of the framework, we run the experiment on different LLMs.
5.3.1 Framework Generalizability
We first test whether our framework can be transfered on different LLMs, including GPT-5-mini, GPT-4.1-nano, Gemini-2.0-flash, Gemini-2.5-flash-lite. The results are shown in Table 3, We can see that all these models achieve better TPR at higher GAN iteration compared with lower GAN iteration, indicating that this framework can be applied on various LLMs. For more detailed results, please check the Appendix A.5.1.
5.3.2 Prompt Transferability
We also investigated the transferability of the generated defense prompts across various large-language models (LLMs). Specifically, we examine whether the prompts optimized using one model (in this case, GPT-4.1-mini) retain their effectiveness when applied to different LLMs (in this case, GPT-4.1-nano, Gemini-2.5-flash, and Gemini-2.5-flash-lite) without modification.
Table 4 illustrates the results for all models at GAN iteration 0, iteration 4, and iteration 8. For stronger LLMs (GPT-4.1-mini, Gemini-2.5-flash), they achieve a high TPR at iteration 8, while weaker LLMs (GPT-4.1-nano, Gemini-2.5-flash-lite) achieve lower TPR at iteration 8. However, we can see that the results for all LLMs improve when more GAN iterations are trained. For more detailed results, please check the Appendix A.5.2.
Gan Iteration GPT-4.1-mini (source) GPT-4.1-nano Gemini-2.5-flash Gemini-2.5-flash-lite 0 0.08 0.01 0.62 0.04 4 0.78 0.07 0.91 0.21 8 0.84 0.15 0.98 0.39
6 Ablation Study
Figure 5: Iterative performance of the attacker and defender across GAN iterations, measured by True Positive Rate (TPR). Shaded region represent the standard deviation across runs. The default method has the steadiest improvement and achieve the best TPR at last.
To understand the contribution of each component in AEGIS , we conducted several ablation studies. These studies involve systematically removing or altering parts of our system and observing the impact on performance.
6.1 Without Gradient Buffer
We removed the gradient buffer, which stores historical textual gradients for the optimization prompt. This change leads to a slower convergence rate and degrades the TPR in defense about 5%.
6.2 Without Multiple Gradients
We simplify the gradient generation process to use only one optimization type. For the defender, we used only the True Positive Rate (TPR), and for the attacker, only the Attack Success Rate (ASR). This led to a performance degrade in both attacker and defender, where the final performance of the defense degrades with over 10
6.3 Single-Sided Training
We also experimented with training only one side of the GAN framework. The lack of an adaptive adversary meant that the trained model quickly overfit to its static opponent, highlighting the importance of co-evolution. When training the defender against a static set of attacks, it fails to generalize to new, unseen attacks, resulting in a lower overall TPR.
The comparisons between all the ablation studies are presented in the figures above. Fig 5 compares the TPR between each ablation, and Fig 4 compares the TNR between each ablation.
7 Conclusion
Through extensive experiments and ablation studies, we confirmed the importance of co-evolution, gradient replay, and multi-objective optimization. These findings suggest that adversarial training, when applied at the prompt level, offers a scalable and effective solution for safeguarding LLMs in real-world deployments. In future work, we plan to extend AEGIS to more complex scenarios and improve prompt interpretability.
Limitation
Despite the promising results, our study has several limitations. First, our evaluation focused on automated assignment grading task, which may not fully capture the diversity of real-world security-sensitive tasks and show the generalizability of our framework. Second, our defense method mainly targets text-based dialogue systems, and its effectiveness in multimodal systems remains unclear. Third, while we focused on quantitative evaluation of attack success rates and defense robustness, large-scale human evaluations were not conducted. Addressing these limitations in future work will be important for building more comprehensive and deployable defense systems.
Appendix A Appendix
A.1 Algorithm for Adversarial Co-evolution Framework
Algorithm 1 outlines the overall adversarial co-evolution procedure. In each iteration, the attacker and defender are alternately optimized through a prompt-based generation and evaluation process. Candidates are produced using the prompt optimization framework $TGO^{+}$, and their effectiveness is evaluated using task-specific criteria by $Eval()$. From these candidates, the best-performing attacker and defender are selected at the end of each iteration via $Val()$. This co-evolutionary process continues until the maximum number of iterations $N$ is reached, at which point the framework outputs the final attacker and defender models, $ATK^{N}_{best}$ and $DEF^{N}_{best}$.